### v1

# In-plane transverse polarization in heavy-ion collisions

Abstract
We give an analytical expression for the last component of the spin polarization $P^{x}$, the in-plane polarization, in heavy-ion collisions that has, to our knowledge, not been discussed in theories nor measured in heavy-ion collision experiments. We also carry out a numerical study of $P^{x}$ using a hydrodynamic model simulation as a cross-check for the analytical formula. It is found that if the temperature-gradient contribution is neglected the simulation result for $P^{x}$ qualitatively agrees with the analytical one. The prediction of $P^{x}$ can be tested in experiments and will contribute to provide a complete and consistent picture of spin phenomena in heavy-ion collisions.
I Introduction
A substantial part of the orbital angular momentum in heavy-ion collisions can be transferred into the strong interaction matter and leads to the spin polarization of final-state particles through spin-orbit coupling (Liang:2004ph, ). This particular type of polarization is parallel to the normal vector of the reaction plane formed by the impact parameter and the beam direction which is fixed for all particles in one single event. As such, it is usually referred to as the global spin polarization to distinguish it from the spin polarization effects in proton-proton (Bunce:1976yb, ; Heller:1978ty, ; Thomas:1978ova, ; Wilkinson:1981jy, ; DeGrand:1980gc, ; DeGrand:1981pe, ; Heller:1996pg, ) and electron-positron collisions (Belle:2018ttu, ). There, the spin polarization of particles is normally with respect to the production plane spanned by the particle’s momentum and the beam direction which vary for particles with different momenta even in the same event. Early studies of the global spin polarization can be found in Refs. (Liang:2004ph, ; Voloshin:2004ha, ; Gao:2007bc, ; Betz:2007kg, ; Becattini:2007sr, ).
The global spin polarization of $\Lambda$ hyperons has been measured across a wide range of collision energies (STAR:2007ccu, ; STAR:2017ckg, ; STAR:2018gyt, ; HADES:2022enx, ; ALICE:2019onw, ). Driven by experimental progress, the global spin polarization has been extensively studied in theoretical models. In Ref. (Liang:2004ph, ), a theoretical model was proposed in which unpolarized quarks are scattered at fixed impact parameters by a static potential, leading to polarized quarks after scatterings as the result of the spin-orbit coupling. The model was later applied to multiple scatterings by the static potential (Huang:2011ru, ). As a major improvement of the model, a formalism for two-to-two quark scatterings at fixed impact parameters was developed (Gao:2007bc, ). These results (Liang:2004ph, ; Gao:2007bc, ) are only for one single scattering. In a thermal system, particle collisions take place with arbitrary incident momenta. Therefore, one has to take ensemble average over all possible collisions in order to obtain an average effect (Zhang:2019xya, ). In this way, it was shown that the spin polarization arising from the spin-orbit coupling in one single scattering can be converted to that from the spin-vorticity coupling after ensemble average over all possible collisions in a thermal system (Zhang:2019xya, ). In practice, spin polarization in hydrodynamical and transport models (Xia:2018tes, ; Karpenko:2016jyx, ; Sun:2017xhx, ; Li:2017slc, ; Wei:2018zfb, ; Vitiuk:2019rfv, ; Becattini:2020ngo, ; Fu:2020oxj, ; Ryu:2021lnx, ; Jiang:2023vxp, ; Palermo:2024tza, ) is calculated by mapping the vorticity to the spin polarization on the freeze-out hypersurface (Becattini:2013fla, ; Fang:2016vpj, ). For recent reviews on global spin polarizations, see Refs. (Wang:2017jpl, ; Florkowski:2018fap, ; Huang:2020dtn, ; Gao:2020lxh, ; Gao:2020vbh, ; Liu:2020ymh, ; Becattini:2022zvf, ; Hidaka:2022dmn, ; Becattini:2024uha, ).
In addition to the global spin polarization, the spin polarization along the beam direction was also proposed in hydrodynamic and transport models (Becattini:2017gcx, ; Xia:2018tes, ) with the expected behavior $P^{z}\sim-\sin(2\phi_{p})$ where $\phi_{p}$ represents the transverse azimuthal angle of the hyperon’s momentum in the reaction plane. However the experimental measurement (STAR:2019erd, ) show an opposite sign behaviour $P^{z}\sim\sin(2\phi_{p})$. The first theoretical explanation of this discrepancy was provided in Ref. (Wu:2019eyi, ; Wu:2020yiz, ), where it was found that the temperature vorticity qualitatively accounts for spin polarization along the beam direction as well as the global spin polarization as a funnction of $\phi_{p}$. Later on, it was found that the contribution from the shear stress tensor can also yield the correct sign for longitudinal polarization (Fu:2021pok, ; Becattini:2021iol, ; Yi:2021ryh, ; Florkowski:2021xvy, ; Wagner:2022gza, ; Wu:2022mkr, ). More recently, it has been proposed that in a thermal model the projected thermal vorticity along with dissipative corrections can describe the behavior of longitudinal polarization (Banerjee:2024xnd, ). Very recently dissipative relativistic spin hydrodynamics was developed from quantum kinetic theory for massive particles with non-local collisions (Weickgenannt:2020aaf, ; Weickgenannt:2021cuo, ; Weickgenannt:2022zxs, ) and gave a good description of longitudinal polarization data (Sapna:2025yss, ). These theoretical approaches collectively indicate that global equilibrium has not been achieved, necessitating the inclusion of off-equilibrium effects in the analysis (Sheng:2021kfc, ).
Surprisingly, there is a simple way to explain the behavior of the longitudinal polarization. If one uses the non-relativistic approximation in the blast-wave model (STAR:2019erd, ; Voloshin:2017kqp, ), one can show that $P^{z}\sim\omega^{xy}\sim(1/r)v_{2}v_{r}\sin(2\phi)$ where the profile of the transverse radial flow velocity is given by $\mathbf{v}\sim\mathbf{e}_{r}v_{r}\left[1+v_{2}\cos(2\phi)\right]$, $\omega^{xy}\sim\nabla_{x}v^{y}-\nabla_{y}v^{x}$ is the longitudinal component of the vorticity vector, and $v_{r}$ and $\mathbf{e}_{r}=(\cos\phi,\sin\phi)$ denote the radial flow velocity and its direction, respectively. This provides a straightforward explanation for the experimentally observed pattern of the longitudinal spin polarization. However, it is not clear whether such a simple non-relativistic approximation (STAR:2019erd, ; Voloshin:2017kqp, ) could describe other spin observables without relativistic effects that could potentially alter the observed pattern in experiments.
Inspired by the simple and intuitive blast wave picture of the longitudinal spin polarization, we had performed a comprehensive analysis of spin observables in the framework of the modified or extended blast wave model (Retiere:2003kf, ; Jaiswal:2015saa, ; Yang:2016rnw, ; Yang:2020oig, ; Yang:2022yxa, ). We found analytical expressions or solutions for the longitudinal and global spin polarizations as functions of particle’s momentum and collision centrality under flow-momentum correspondence (Arslan:2024dwi, ). From the analytic solutions, one could clearly see that the global spin polarization is driven by the directed flow, while the longitudinal spin polarization is driven by the ellipticity in flow. The analytical solutions can be improved systematically by perturbative expansion in the small deviations from the flow-momentum correspondence.
In this paper, we provide an analytical expression for the last component of the spin vector $P^{x}$, the in-plane spin polarization, that has, to our knowledge, not been discussed in theories nor measured in experiments. We also carry out a numerical study of $P^{x}$ using a hydrodynamic simulation as a cross-check for such an analytical solution. The prediction of $P^{x}$ can be tested in experiments and may provide a complete and consistent picture for spin phenomena in heavy-ion collisions.
The paper is organized as follows. In Sec. II we introduce the approach to calculating the spin polarization observables. In Sec. III, we describe the perturbation method based on the flow-momentum correspondence. In Sec. IV, we present the analytical expressions for the in-plane spin polarization $P^{x}$, the main results of the paper. Section VI presents numerical results for analytical solutions and hydrodynamic simulations. A summary of the paper is given in the final section.
II Model description
The extended blast wave model provides a unified framework for description of transverse mass spectra, elliptic flow, and two-particle correlations (Retiere:2003kf, ; Jaiswal:2015saa, ; Yang:2016rnw, ; Yang:2020oig, ; Yang:2022yxa, ). It offers a straightforward parameterization of the system at kinetic freeze-out, characterized by the temperature, transverse flow, and transverse radius of the source (Siemens:1978pb, ; Lee:1990sk, ; Schnedermann:1993ws, ; Huovinen:2001cy, ).
We consider the non-central collision of two high-energy nuclei moving with the speed of light along the $\pm z$ direction at $x=\pm b/2$. The direction perpendicular to the reaction plane is then the $y$ direction.
The flow four-velocity and the particle’s four-momentum can be parameterized as
$u^{\mu}(x)=$ $\left(\cosh\eta\cosh\rho,\sinh\rho\cos\phi_{b},\sinh\rho\sin\phi_{b},\sinh\eta\cosh\rho\right),$ (1)
$p^{\mu}=$ $\left(m_{T}\cosh Y,p_{T}\cos\phi_{p},p_{T}\sin\phi_{p},m_{T}\sinh Y\right).$ (2)
Here $\eta$ and $Y$ are the space-time and momentum rapidity respectively, $p_{T}$ is the transverse momentum, $m_{T}=\sqrt{m^{2}+p_{T}^{2}}$ is the transverse mass, $\phi_{b}$ is the azimuthal angle in the transverse flow, the transverse expansion of the fireball (Lee:1990sk, ; Huovinen:2001cy, ; Retiere:2003kf, ) is described by the transverse rapidity $\rho$ as a function of $r$ (transverse radius), $\phi_{b}$ (azimuthal angle of the flow velocity in transverse plane), $\epsilon$ (the eccentricity parameter in transverse emission area), and $\eta$ as follows
$\rho\left(r,\phi_{b},\eta\right)\approx\frac{r}{R}\left[\rho_{0}+\rho_{1}(\eta)\cos(\phi_{b})+\left(\rho_{2}+\frac{1}{2}\epsilon\rho_{0}\right)\cos(2\phi_{b})\right],$ (3)
where $\rho_{0}$ characterizes the mean transverse rapidity of the source element, $\rho_{1}(\eta)=\alpha_{1}\eta$ and $\rho_{2}$ describe the azimuthal anisotropy of the transverse rapidity. The transverse emission area can be characterized by $R_{x}$ and $R_{y}$, effective radii of the elliptic source in $x$- and $y$-direction respectively, with $R=(R_{x}+R_{y})/2$. In deriving Eq. (3) we used the approximation $\epsilon\equiv(R_{y}-R_{x})/R\ll 1$ which works very well in describing data (Retiere:2003kf, ). We assume $\alpha_{1}\sim\rho_{2}\sim\epsilon\ll\rho_{0}$, so that $\alpha_{1}$, $\rho_{2}$ and $\epsilon$ can be treated as perturbations relative to $\rho_{0}$. Note that $\phi_{b}$ in the transverse flow velocity is related to the azimuthal angle $\phi_{s}$ in transverse coordinate space of the emission source by $\tan\phi_{b}=(R_{x}^{2}/R_{y}^{2})\tan\phi_{s}$, hence the difference between $\phi_{b}$ and $\phi_{s}$ is $O(\epsilon)$.
The particle’s distribution function in phase space is assumed to follow the Boltzmann distribution, $f(x,p)\equiv f(p\cdot u)=\exp(-\beta p\cdot u)$, where $\beta=1/T$ is the inverse temperature and $p\cdot u$ is given by
$p\cdot u=m_{T}\cosh\rho\cosh(\eta-Y)-p_{T}\sinh\rho\cos(\phi_{b}-\phi_{p}).$ (4)
We see that $f(p\cdot u)$ depends on rapidities and azimuthal angles through $\eta-Y$ and $\phi_{b}-\phi_{p}$. This distribution reaches a maximum when $\eta\approx Y$ and $\phi_{b}\approx\phi_{p}$, i.e. the spacetime and momentum rapidities are equal and the flow and momentum azimuthal angles are equal. If we set $\eta=Y$ and $\phi_{b}=\phi_{p}$, the distribution reaches a maximum at $p_{T}/m_{T}=\tanh\rho$, meaning that the transverse momentum rapidity is equal to the transverse flow rapidity. These equalities between flow variables and particle momentum variables are called the flow-momentum correspondence in the fireball’s expansion.
The observables can be calculated on the freeze-out hypersurface $\Sigma$ defined by equal temperature condition at the freeze-out proper time $\tau=\tau_{f}$, $T(\tau_{f},\eta,r,\phi_{s})=T_{f}$, which is a three-dimensional manifold in $\eta$, $r$ and $\phi_{s}$. On the freeze-out hypersurface, the partial derivatives of $T$ with respect to $\eta$, $r$ and $\phi_{s}$ are assumed to vanish except that with respect to $\tau$, $\left.\partial T/\partial\tau\right|_{\Sigma}\neq 0$. The emission function $S(x,p)$ represents the probability of emitting a particle with the momentum $p$ at the space-time $x$ incorporated with the freeze-out condition (Arslan:2024dwi, ),
$S(x,p)=m_{T}\cosh(\eta-Y)\delta(\tau-\tau_{f})\Theta(R-r)f(p\cdot u).$ (5)
Then the expectation value of an observable as a function of three-momentum (or $p_{T}$, $\phi_{p}$ and $Y$) can be calculated as
$\left\langle O\right\rangle(\mathbf{p})=$ $\frac{\int d^{4}x\hat{O}(x,p)S(x,p)}{\int d^{4}xS(x,p)}$
$=$ $\frac{\int_{0}^{R}dr\int_{-\infty}^{\infty}d\eta\int_{0}^{2\pi}d\phi_{s}\;rm_{T}\cosh(\eta-Y)\exp(-\beta p\cdot u)\hat{O}(x,p)}{\int_{0}^{R}dr\int_{-\infty}^{\infty}d\eta\int_{0}^{2\pi}d\phi_{s}\;rm_{T}\cosh(\eta-Y)\exp(-\beta p\cdot u)},$ (6)
where $\hat{O}(x,p)$ is the observable in phase space and $d^{4}x=\tau rd\tau d\eta drd\phi_{s}$. The observable’s spectra in some variables of $p_{T}$, $\phi_{p}$ and $Y$ can be obtained by integration over the rest of the variables in both numerator and denominator in Eq. (6).
III Perturbation method and flow-momentum correspondence
From Eqs. (3,4,5), the distribution function $f(p\cdot u)$ depends on $\phi_{b}$, so we have to convert the integral over $\phi_{s}$ in Eq. (6) to that over $\phi_{b}$ by rewriting $\phi_{s}$ in $\hat{O}(x,p)$ in terms of $\phi_{b}$. The integrals over $\eta$ and $\phi_{s}$ (or equivalently $\phi_{b}$) can be carried out using a perturbation method. To $O(\epsilon)$, the Boltzmann distribution function can be approximated as
$\exp(-\beta p\cdot u)\approx$ $\exp\left[-\beta m_{T}\cosh\bar{\rho}\cosh(\Delta\eta)+\beta p_{T}\sinh\bar{\rho}\cos(\Delta\phi)\right]$
$\times\left[1-\delta\rho\beta m_{T}\sinh\bar{\rho}\cosh(\Delta\eta)+\delta\rho\beta p_{T}\cosh\bar{\rho}\cos(\Delta\phi)\right],$ (7)
where $\bar{\rho}\equiv(r/R)\rho_{0}$, $\Delta\eta\equiv\eta-Y$, $\Delta\phi\equiv\phi_{b}-\phi_{p}$ and $\delta\rho$ is given by
$\delta\rho=$ $\frac{r}{R}\left[\rho_{1}(\eta)\cos(\phi_{b})+\left(\rho_{2}+\frac{1}{2}\epsilon\rho_{0}\right)\cos(2\phi_{b})\right]$
$=$ $\frac{r}{R}\left[\rho_{1}(\Delta\eta+Y)\cos(\Delta\phi)\cos(\phi_{p})-\rho_{1}(\Delta\eta+Y)\sin(\Delta\phi)\sin(\phi_{p})\right$.
$\left.+\left(\rho_{2}+\frac{1}{2}\epsilon\rho_{0}\right)\cos(2\Delta\phi)\cos(2\phi_{p})-\left(\rho_{2}+\frac{1}{2}\epsilon\rho_{0}\right)\sin(2\Delta\phi)\sin(2\phi_{p})\right],$ (8)
We see that $\delta\rho$ depends on $\Delta\eta$, $Y$, $\Delta\phi$ and $\phi_{p}$ and is of $O(\epsilon)$. The integral measure for $\phi_{s}$ in Eq. (6) can be converted to that for $\phi_{b}$ to $O(\epsilon)$ as
$d\phi_{s}\approx$ $d\phi_{b}\left[1+2\epsilon\cos(2\phi_{b})\right]$
$=$ $d\phi_{b}\left[1+2\epsilon\cos(2\Delta\phi)\cos(2\phi_{p})-2\epsilon\sin(2\Delta\phi)\sin(2\phi_{p})\right].$ (9)
We note that the expansion in Eq. (7) is only valid inside the integral since for large $\bar{\rho}$ and $\Delta\eta$, the $\delta\rho$ terms can easily become larger than 1. When integrated over $r$ and $\Delta\eta$, the exponential factor suppresses the contribution from large $\Delta\eta$ ($\cosh\bar{\rho}$ and $\sinh\bar{\rho}$ are finite since $\bar{\rho}<1$ for $\rho_{0}\sim 1$).
We now look at the polarization observables $\hat{O}(x,p)=\hat{P}^{i}(x,p)$ with $i=x,y,z$. Normally they depend on $\eta$, $\phi_{s}$ and $\phi_{b}$ through functions of $\sinh\eta$, $\cosh\eta$, $\sin\phi_{s/b}$ and $\cos\phi_{s/b}$. With $\eta=\Delta\eta+Y$ and $\phi_{b}=\Delta\phi+\phi_{p}$, the polarization observables can be expressed as functions of $\sinh\Delta\eta$, $\cosh\Delta\eta$, $\sin\Delta\phi$ and $\cos\Delta\phi$. The integrals over $\eta$ and $\phi_{s}$ in the numerator in Eq. (6) can be schematically written as
$I_{\eta,\phi}=$ $\int_{-\infty}^{\infty}d\Delta\eta\int_{0}^{2\pi}d\Delta\phi\;\cosh(\Delta\eta)F\left(r,\Delta\eta,\Delta\phi,p_{T},Y,\phi_{p}\right)$
$\times\exp\left[-\beta m_{T}\cosh\bar{\rho}\cosh(\Delta\eta)+\beta p_{T}\sinh\bar{\rho}\cos(\Delta\phi)\right],$ (10)
where the integrand function $F$ is defined as
$F\left(r,\Delta\eta,\Delta\phi,p_{T},Y,\phi_{p}\right)\approx$ $\hat{O}\left(r,\Delta\eta,\Delta\phi,p_{T},Y,\phi_{p}\right)$
$\times\left[1+2\epsilon\cos(2\Delta\phi)\cos(2\phi_{p})-2\epsilon\sin(2\Delta\phi)\sin(2\phi_{p})\right$.
$\left.-\delta\rho\beta m_{T}\sinh\bar{\rho}\cosh(\Delta\eta)+\delta\rho\beta p_{T}\cosh\bar{\rho}\cos(\Delta\phi)\right].$ (11)
The factor inside the square brackets is the product of the second factor in Eq. (7) and the factor in Eq. (9). Note that in Eq. (10) $\bar{\rho}$ only depends on $r$, so the integral $I_{\eta,\phi}$ can be completed through formulas for modified Bessel functions of the first and second kinds.
The flow-momentum correspondence in central rapidity means setting $\eta=Y=0$ in $F\left(r,\Delta\eta,\Delta\phi,p_{T},Y,\phi_{p}\right)$, so $F$ becomes a function of $r$, $p_{T}$, $Y$, $\phi_{s}$($\phi_{b}$) and $\phi_{p}$,
$F\left(r,0,\Delta\phi,p_{T},0,\phi_{p}\right)\approx$ $\hat{O}\left(r,0,\Delta\phi,p_{T},0,\phi_{p}\right)$
$\times\left[1+2\epsilon\cos(2\Delta\phi)\cos(2\phi_{p})-2\epsilon\sin(2\Delta\phi)\sin(2\phi_{p})\right$.
$\left.-\left.\delta\rho\right|_{\eta=Y=0}\beta m_{T}\sinh\bar{\rho}+\left.\delta\rho\right|_{\eta=Y=0}\beta p_{T}\cosh\bar{\rho}\cos(\Delta\phi)\right],$ (12)
where $\left.\delta\rho\right|_{\eta=Y=0}$ is given by
$\left.\delta\rho\right|_{\eta=Y=0}$ $=\frac{r}{R}\left[\left(\rho_{2}+\frac{1}{2}\epsilon\rho_{0}\right)\cos(2\Delta\phi)\cos(2\phi_{p})\right$.
$\left.-\left(\rho_{2}+\frac{1}{2}\epsilon\rho_{0}\right)\sin(2\Delta\phi)\sin(2\phi_{p})\right].$ (13)
Equation (12) is a good approximation at RHIC energy with $\beta p_{T}\delta\rho\ll 1$ being well satisfied but it is not at LHC energy. In this paper we will calculate the expectation values of observables to $O(\epsilon)$ with the flow-momentum correspondence in central rapidity $\eta=Y=0$ but without imposing that in the azimuthal angle, i.e. $\Delta\phi\neq 0$. This is different from Ref. (Arslan:2024dwi, ) in which both $\eta=Y=0$ and $\Delta\phi=0$ were imposed in $F\left(r,\Delta\eta,\Delta\phi,p_{T},Y,\phi_{p}\right)$.
IV Polarization vectors: general results
In this work, we focus on spin-1/2 particles and include polarization only from the thermal vorticity and thermal shear stress tensors. There may be other sources of polarizations (Sheng:2021kfc, ; Banerjee:2024xnd, ). The spin vectors from these two sources are defined as
$\hat{P}_{\omega}^{\mu}=$ $-\frac{1}{4m}\epsilon^{\mu\nu\sigma\tau}(1-f)\omega_{\nu\sigma}p_{\tau},$ (14)
$\hat{P}_{\xi}^{\mu}=$ $-\frac{1}{2m}\epsilon^{\mu\nu\sigma\tau}(1-f)\frac{p_{\tau}p^{\rho}}{E_{p}}\hat{t}_{\nu}\xi_{\rho\sigma},$ (15)
where the vector $\hat{t}$ is chosen to be $\hat{t}^{\mu}=(1,0,0,0)$ corresponding to the normal direction of the freeze-out hyper-surface for $\eta=Y=0$, $\omega^{\mu\nu}$ and $\xi^{\mu\nu}$ denote the thermal vorticity and thermal shear stress tensors respectively defined as
$\omega^{\mu\nu}=$ $-\frac{1}{2}\left[\partial^{\mu}\left(\beta u^{\nu}\right)-\partial^{\nu}\left(\beta u^{\mu}\right)\right]$
$=$ $-\frac{1}{2T}\left(\partial^{\mu}u^{\nu}-\partial^{\nu}u^{\mu}\right)+\frac{1}{2T^{2}}\left(u^{\nu}\partial^{\mu}T-u^{\mu}\partial^{\nu}T\right),$ (16)
$\xi^{\mu\nu}=$ $\frac{1}{2}\left[\partial^{\mu}\left(\beta u^{\nu}\right)+\partial^{\nu}\left(\beta u^{\mu}\right)\right]$
$=$ $\frac{1}{2T}\left(\partial^{\mu}u^{\nu}+\partial^{\nu}u^{\mu}\right)-\frac{1}{2T^{2}}\left(u^{\nu}\partial^{\mu}T+u^{\mu}\partial^{\nu}T\right).$ (17)
We see that both $\omega^{\mu\nu}$ and $\xi^{\mu\nu}$ can be decomposed into kinetic and T-gradient parts. Note that the definition in Eq. (14) is in a Lorentz covariant form and includes the relativistic effect, different from $\hat{P}_{\omega}^{z}\sim\omega^{xy}$ that was used in Ref. (STAR:2019erd, ).
With the flow velocity given by Eq. (1), the thermal vorticity and thermal shear stress tensors can be evaluated and give $\hat{P}_{\omega}^{x,y,z}$ as functions of $(\eta,r,\phi_{s},\phi_{b})$ and $(Y,p_{T},\phi_{p})$. The expectation values of polarization vectors as functions of $\phi_{p}$ can be obtained from Eq. (6) by integration over $Y$ and $p_{T}$ in both the numerator and denominator as
$P^{x}(\phi_{p})=$ $\left\langle\hat{P}^{x}\right\rangle(\phi_{p})=-\alpha_{1}\frac{1}{16mT\tau R}\sin(2\phi_{p})$
$\times\frac{1}{N_{0}}\left[N_{c}(0|2,3,0)+N_{c}(2|2,3,0)+2N_{c}(2|2,1,2)-4N_{s}(1|2,2,1)\right],$
$P^{y}(\phi_{p})=$ $\left\langle\hat{P}^{y}\right\rangle(\phi_{p})$
$=$ $\alpha_{1}\frac{1}{8mT\tau R}\frac{1}{N_{0}}\left[N_{c}(0|2,1,2)-N_{c}(2|2,1,2)\right]+\alpha_{1}\frac{1}{8mT\tau R}\cos^{2}\phi_{p}$
$\times\frac{1}{N_{0}}\left[N_{c}(0|2,3,0)+N_{c}(2|2,3,0)+2N_{c}(2|2,1,2)-4N_{s}(1|2,2,1)\right],$
$P^{z}(\phi_{p})=$ $\left\langle\hat{P}^{z}\right\rangle(\phi_{p})$
$=$ $\frac{1}{8mT}\sin(2\phi_{p})\frac{1}{N_{0}}\int_{0}^{p_{T}^{\mathrm{max}}}dp_{T}\int_{0}^{R}dr\;rp_{T}$
$\times\sum_{n=0}\left(m_{T}C_{z,n}^{\omega}+p_{T}C_{z,n}^{\xi}\right)K_{1}(\beta m_{T}\cosh\bar{\rho})I_{n}(\beta p_{T}\sinh\bar{\rho}),$ (18)
where $\hat{P}^{i}=\hat{P}_{\omega}^{i}+\hat{P}_{\xi}^{i}$ with $i=x,y,z$ and
$N_{0}=$ $\int_{0}^{p_{T}^{\mathrm{max}}}dp_{T}\int_{0}^{R}dr\;rp_{T}m_{T}\;K_{1}(\beta m_{T}\cosh\bar{\rho})I_{0}(\beta p_{T}\sinh\bar{\rho}),$
$N_{c}(n|n_{1},n_{2},n_{3})=$ $\int_{0}^{p_{T}^{\mathrm{max}}}dp_{T}\int_{0}^{R}dr\;r^{n_{1}}p_{T}^{n_{2}}m_{T}^{n_{3}}\;\cosh\bar{\rho}\;K_{1}(\beta m_{T}\cosh\bar{\rho})I_{n}(\beta p_{T}\sinh\bar{\rho}),$
$N_{s}(n|n_{1},n_{2},n_{3})=$ $\int_{0}^{p_{T}^{\mathrm{max}}}dp_{T}\int_{0}^{R}dr\;r^{n_{1}}p_{T}^{n_{2}}m_{T}^{n_{3}}\;\sinh\bar{\rho}\;K_{1}(\beta m_{T}\cosh\bar{\rho})I_{n}(\beta p_{T}\sinh\bar{\rho}).$ (19)
Similarly the integrated polarization vectors can be obtained by further integrating over $\phi_{p}$ from Eq. (18). But $P^{x}(\phi_{p})$ and $P^{z}(\phi_{p})$ are proportional to $\sin(2\phi_{p})$ whose integration over $\phi_{p}$ is vanishing. In order to obtain meaningful results for integrated polarization vectors in $x$ and $z$ directions, we can calculate weighted observables $P_{\sin 2\phi}^{x}=\left\langle\hat{P}^{x}\sin(2\phi_{p})\right\rangle$ and $P_{\sin 2\phi}^{z}=\left\langle\hat{P}^{z}\sin(2\phi_{p})\right\rangle$. For the longitudinal polarization, we can calculate $P^{y}=\left\langle\hat{P}^{y}\right\rangle$ or $P_{\cos 2\phi}^{y}=\left\langle\hat{P}^{y}\cos(2\phi_{p})\right\rangle$. The results are
$P_{\sin 2\phi}^{x}=$ $\left\langle\hat{P}^{x}\sin(2\phi_{p})\right\rangle=-\alpha_{1}\frac{1}{32mT\tau R}$
$\times\frac{1}{N_{0}}\left[N_{c}(0|2,3,0)+N_{c}(2|2,3,0)+2N_{c}(2|2,1,2)-4N_{s}(1|2,2,1)\right],$
$P^{y}=$ $\left\langle\hat{P}^{y}\right\rangle=\alpha_{1}\frac{1}{16mT\tau R}$
$\times\frac{1}{N_{0}}\left[N_{c}(0|2,3,0)+N_{c}(2|2,3,0)+2N_{c}(0|2,1,2)-4N_{s}(1|2,2,1)\right],$
$P_{\cos 2\phi}^{y}=$ $\left\langle\hat{P}^{y}\cos(2\phi_{p})\right\rangle=\alpha_{1}\frac{1}{32mT\tau R}$
$\times\frac{1}{N_{0}}\left[N_{c}(0|2,3,0)+N_{c}(2|2,3,0)+2N_{c}(2|2,1,2)-4N_{s}(1|2,2,1)\right],$
$P_{\sin 2\phi}^{z}=$ $\left\langle\hat{P}^{z}\sin(2\phi_{p})\right\rangle$
$=$ $\frac{1}{16mT}\frac{1}{N_{0}}\int_{0}^{p_{T}^{\mathrm{max}}}dp_{T}\int_{0}^{R}dr\;rp_{T}$
$\times\sum_{n=0}\left(m_{T}C_{z,n}^{\omega}+p_{T}C_{z,n}^{\xi}\right)K_{1}(\beta m_{T}\cosh\bar{\rho})I_{n}(\beta p_{T}\sinh\bar{\rho}).$ (20)
The only difference between $P^{y}$ and $P_{\cos 2\phi}^{y}$ is that the former has a term $2N_{c}(0|2,1,2)$ while the latter has a term $2N_{c}(2|2,1,2)$. Apart from such a difference, we observe an approximated equality for $P^{y}$ and $P_{\sin 2\phi}^{x}$: $P^{y}\approx-2P_{\sin 2\phi}^{x}$, which will be confirmed by numerical results in Figs. 5, 6 and 7. Furthermore, there is an exact equality for $P_{\sin 2\phi}^{x}$ and $P_{\cos 2\phi}^{y}$: $P_{\sin 2\phi}^{x}=-P_{\cos 2\phi}^{y}$, showing a good symmetry between the in-plane and out-of-plane polarization.
In a similar way, the expectation values of polarization or weighted polarization vectors as functions of $p_{T}$ can be obtained from Eq. (6) by integration over $Y$ and $\phi_{p}$ in both the numerator and denominator as
$P_{\sin 2\phi}^{x}(p_{T})=$ $\left\langle\hat{P}^{x}\sin(2\phi_{p})\right\rangle(p_{T})=-\alpha_{1}\frac{1}{32mT\tau R}$
$\times\frac{1}{N_{0}(p_{T})}\left[N_{c}^{p}(0|2,2,0)+N_{c}^{p}(2|2,2,0)+2N_{c}^{p}(2|2,0,2)-4N_{s}^{p}(1|2,1,1)\right],$
$P^{y}(p_{T})=$ $\left\langle\hat{P}^{y}\right\rangle(p_{T})=\alpha_{1}\frac{1}{16mT\tau R}$
$\times\frac{1}{N_{0}(p_{T})}\left[N_{c}^{p}(0|2,2,0)+N_{c}^{p}(2|2,2,0)+2N_{c}^{p}(0|2,0,2)-4N_{s}^{p}(1|2,1,1)\right],$
$P_{\cos 2\phi}^{y}(p_{T})=$ $\left\langle\hat{P}^{y}\cos(2\phi_{p})\right\rangle(p_{T})=\alpha_{1}\frac{1}{32mT\tau R}\frac{1}{N_{0}(p_{T})}$
$\times\left[N_{c}^{p}(0|2,2,0)+N_{c}^{p}(2|2,2,0)+2N_{c}^{p}(2|2,0,2)-4N_{s}^{p}(1|2,1,1)\right],$
$P_{\sin 2\phi}^{z}(p_{T})=$ $\left\langle\hat{P}^{z}\sin(2\phi_{p})\right\rangle(p_{T})$
$=$ $\frac{1}{16mT}\frac{1}{N_{0}(p_{T})}\int_{0}^{R}dr\;r\sum_{n=0}\left(m_{T}C_{z,n}^{\omega}+p_{T}C_{z,n}^{\xi}\right)$
$\times K_{1}(\beta m_{T}\cosh\bar{\rho})I_{n}(\beta p_{T}\sinh\bar{\rho}),$ (21)
where $N_{0}(p_{T})$ and $N_{i}^{p}(n|n_{1},n_{2},n_{3})$ ($i=c,s$) are functions of $p_{T}$ defined as
$N_{0}(p_{T})=$ $\int_{0}^{R}dr\;rm_{T}\;K_{1}(\beta m_{T}\cosh\bar{\rho})I_{0}(\beta p_{T}\sinh\bar{\rho}),$
$N_{c}^{p}(n|n_{1},n_{2},n_{3})=$ $\int_{0}^{R}dr\;r^{n_{1}}p_{T}^{n_{2}}m_{T}^{n_{3}}\;\cosh\bar{\rho}\;K_{1}(\beta m_{T}\cosh\bar{\rho})I_{n}(\beta p_{T}\sinh\bar{\rho}),$
$N_{s}^{p}(n|n_{1},n_{2},n_{3})=$ $\int_{0}^{R}dr\;r^{n_{1}}p_{T}^{n_{2}}m_{T}^{n_{3}}\;\sinh\bar{\rho}\;K_{1}(\beta m_{T}\cosh\bar{\rho})I_{n}(\beta p_{T}\sinh\bar{\rho}).$ (22)
We observe that $N_{0}$ and $N_{i}(n|n_{1},n_{2},n_{3})$ ($i=c,s$) in (19) can be obtained by integration of $N_{0}(p_{T})$ and $N_{i}^{p}(n|n_{1},n_{2},n_{3})$ over $p_{T}$ as
$N_{0}=$ $\int_{0}^{p_{T}^{\mathrm{max}}}dp_{T}p_{T}N_{0}(p_{T}),$
$N_{i}(n|n_{1},n_{2}+1,n_{3})=$ $\int_{0}^{p_{T}^{\mathrm{max}}}dp_{T}p_{T}N_{i}^{p}(n|n_{1},n_{2},n_{3}).$ (23)
We also observe an approximated equality for $P^{y}(p_{T})$ and $P_{\sin 2\phi}^{x}(p_{T})$ from Eq. (21): $P^{y}\approx-2P_{\sin 2\phi}^{x}$, which will be confirmed by numerical results in Figs. 5, 6 and 7. There is also an exact equality for $P_{\sin 2\phi}^{x}(p_{T})$ and $P_{\cos 2\phi}^{y}(p_{T})$: $P_{\sin 2\phi}^{x}(p_{T})=-P_{\cos 2\phi}^{y}(p_{T})$, which shows a good symmetry between the in-plane and out-of-plane polarization.
V Transverse polarizations
We see from Eqs. (18,20,21) that both $P^{x}$ and $P^{y}$ are driven by the directed flow. The $\phi_{p}$ pattern of $P^{x}$ is $\sin(2\phi_{p})$, same as $P^{z}$ but with a slightly smaller magnitude which will be shown in Fig. 7. There are two terms in $P^{y}(\phi_{p})$, the first term would be vanishing if we take $\Delta\phi=0$ in $\hat{P}^{y}(\phi_{p})$ before the integration over $\Delta\phi$ in calculating its expectation value as in Ref. (Arslan:2024dwi, ), while the second term is proportional to $P^{x}(\phi_{p})$ up to a modulation factor $\sim\sin\phi_{p}$. We can combine $P^{x}(\phi_{p})$ and $P^{y}(\phi_{p})$ to form a vector in the transverse plane
$\mathbf{P}_{T}(\phi_{p})=$ $\mathbf{e}_{x}P^{x}(\phi_{p})+\mathbf{e}_{y}P^{y}(\phi_{p})$
$=$ $\mathbf{e}_{\phi}\alpha_{1}\frac{1}{8mT\tau R}\cos\phi_{p}\frac{1}{N_{0}}\left[N_{c}(0|2,3,0)+N_{c}(2|2,3,0)\right$.
$\left.+N_{c}(0|2,1,2)+N_{c}(2|2,1,2)-4N_{s}(1|2,2,1)\right]$
$+\mathbf{e}_{r}\alpha_{1}\frac{1}{8mT\tau R}\sin\phi_{p}\frac{1}{N_{0}}\left[N_{c}(0|2,1,2)-N_{c}(2|2,1,2)\right]$ (24)
where we have used $\mathbf{e}_{\phi}=-\mathbf{e}_{x}\sin\phi_{p}+\mathbf{e}_{y}\cos\phi_{p}$ and $\mathbf{e}_{y}=\mathbf{e}_{r}\sin\phi_{p}+\mathbf{e}_{\phi}\cos\phi_{p}$. The geometric configuration for $\mathbf{P}_{T}(\phi_{p})$ is shown in Fig. 1. The vector of $\mathbf{P}_{T}(\phi_{p})$ with arrow and length is plotted in Fig. 2 along an ellipse varying with $\phi_{p}$ (note that $\alpha_{1}<0$).
Figure 1: The transverse spin polarization vector $\mathbf{P}_{T}(\phi_{p})$ in Eq. (24).
Figure 2: Vector plot of $\mathbf{P}_{T}(\phi_{p})$ with arrow and length on a circle varying with $\phi_{p}$.
VI Numerical results
In this section, we will give numerical results for polarization observables using analytical formulas derived in Sec. IV. A systematic comparison between numerical results for $P^{y}$ and $P^{z}$ with experimental data will be performed. We will also give numerical results for the in-plane polarization $P^{x}$ by hydrodynamical simulation and compare them with the prediction based on analytical formulas.
VI.1 Numerical results from analytical formulas
The parameters we choose for Au+Au collisions at $\sqrt{s_{NN}}=200$ GeV and different centralities are listed in Table 1. The main difference in the values of parameters between this paper and Ref. (Arslan:2024dwi, ) is in $\rho_{2}$ and $\epsilon$ which are determined by fitting the directed and elliptic flow data. In Ref. (Arslan:2024dwi, ), different combinations of $\rho_{2}$ and $\epsilon$ can fit the flow data equally well, but the new values of $\rho_{2}$ and $\epsilon$ in this paper are better suited for the current analytical formulation of the polarization. In Fig. 3 we show the model fit to elliptic flow data of light particles in 10-80% central Au+Au collisions with the values of $\rho_{2}$ and $\epsilon$ listed in Table 1. The freeze-out temperature $T_{f}$ and transverse rapidity parameter $\rho_{0}$ are extracted by fitting transverse momentum spectra (STAR:2004jwm, ; STAR:2008med, ). The parameter $\alpha_{1}$ is set to 0.05-0.05 by fitting the data for directed flows of $\Lambda/\overline{\Lambda}$ in 10-40% central Au+Au collisions (STAR:2017okv, ). For elliptic flows, we fit the data for light particles and $\Lambda+\overline{\Lambda}$ in different centralities (STAR:2004jwm, ; STAR:2008ftz, ). Our results for light particles in 10-80% central collisions (STAR:2004jwm, ) and $\Lambda+\overline{\Lambda}$ in 10-40% central collisions (STAR:2008ftz, ) agree well with experimental data.
centrality $R$ (fm) $T$ (MeV) $\rho_{0}$ $\rho_{2}$ $\epsilon$ $\alpha_{1}$ $\tau_{f}$(fm/c)
10%-20% 11.5 99.5 0.982 0.043 0.043 -0.05 7.8
20%-30% 10.3 102 0.937 0.054 0.055 -0.05 6.9
30%-40% 9 104 0.894 0.072 0.062 -0.05 5.1
40%-50% 7.8 107 0.841 0.08 0.068 -0.05 3.3
50%-60% 7 110 0.788 0.092 0.073 -0.05 2.6
60%-70% 6.3 116 0.707 0.102 0.078 -0.05 2.3
70%-80% 5.5 125 0.608 0.122 0.082 -0.05 2
Table 1: Parameters used in this paper for Au+Au collisions at $\sqrt{s_{NN}}=200$ GeV.
The spin polarizations $P^{z}$ and $P_{H}\equiv-P^{y}$ as functions of $\phi_{p}$ are calculated by Eq. (18). The experimental data for $P^{z}$ are available for the 20-60% centrality Au+Au collisions at $\sqrt{s_{NN}}=200\text{ GeV}$ (STAR:2019erd, ), while the $P^{y}$ data are available for the 20-50% centrality class (STAR:2018gyt, ). The comparison between the calculated results and experimental data are shown in Fig. 4. These results are different from Ref. (Voloshin:2017kqp, ) in which a non-relativistic approximation $P^{z}\approx\omega^{z}/2$ was used.
Figure 3: The model fit to elliptic flow data of light particles in 10-80% central Au+Au collisions at 200 GeV with the values of $\rho_{2}$ and $\epsilon$ listed in Table 1.
Figure 4: The results for $P^{z}$ (right panel) and $P_{H}\equiv-P^{y}$ (left panel) as functions of $\phi_{p}$ following Eq.(20). The black solid lines represent the calculated results based on analytical formulas.
Figure 5: The calculated results for $\left\langle P^{z}\sin(2\phi_{p})\right\rangle$ and $P_{H}$ as functions of $p_{T}$ from Eq. (21). The polarization at 20-60% and 20-50% is calculated by the total particle-production weighted average.
Figure 6: The centrality dependence of $P_{z}$ and $P_{H}$.
The calculated results for the transverse momentum dependence of polarization following the analytical formulas in Eq. (21). The results are shown in Fig. 5. The data show that $P_{\sin 2\phi}^{z}(p_{T})$ is almost a constant when $p_{T}\gtrsim 1\text{ GeV}$ and this behavior can be well described by the analytical formula in Eq. (21). In contrast, the result based on the analytical formula in Ref. (Arslan:2024dwi, ) grows with increasing $p_{T}$. The improved results obtained in this work therefore indicate that relativistic formulation is important in describing $P_{\sin 2\phi}^{z}(p_{T})$ correctly.
We can also calculate the centrality dependence of $P^{z}$ in the form of $\left\langle\hat{P}^{z}\sin(2\phi_{p})\right\rangle$ and $P_{H}$ using the analytical formulas in Eq. (20). The comparison of theoretical results with data is shown in Fig. 6. One can see that theoretical curves grow from central to peripheral collisions which can describe the experimental data.
With the values of parameters in Table 1 fixed by experimental data for collective flows and polarizations along the longitudinal and global orbital angular momentum directions, $P^{z}$ and $P^{y}$, we can make prediction for a new observable, the in-plane polarization $P^{x}$. The numerical results from analytical formula are presented in Fig. 7. We see that the behavior and magnitude of $P^{x}$ are similar to $P^{z}$ although the former is driven by the directed flow while the latter is driven by the elliptic flow.
Figure 7: The first panel: the in-plane polarization $P^{x}$ as functions of $\phi_{p}$; The second and third panels: $\left\langle\hat{P}^{x}\sin(2\phi_{p})\right\rangle$ as functions of $p_{T}$ and centrality.
VI.2 Simulation results by hydrodynamical models
To compare and understand the polarization results from the blast wave model, we also calculate the corresponding polarization observables using the realistic (3+1)D iEBE-MUSIC framework Schenke:2010nt ; Schenke:2010rr ; Paquet:2015lta ; Denicol:2018wdp . This framework includes smooth initial conditions from SMASH, realistic viscous hydrodynamic evolution, the iSS sampler, and the SMASH afterburner. Specifically, for smooth initial conditions from SMASH SMASH:2016zqf ; Schafer:2021csj , we use a Gaussian smearing function to construct the initial energy-momentum tensor $T_{0}^{\mu\nu}$ and net baryon current $J_{0}^{\mu}$ when the initial hadrons reach the hypersurface at the initial proper time $\tau_{0}=$0.5 fm. The Gaussian widths are $\sigma_{r}=$1.0 fm and $\sigma_{\eta_{s}}=$0.8. Each smooth initial condition is generated by averaging over 5000 events in each 10% centrality bin.
The subsequent hydrodynamic evolution included the net baryon current without diffusion. The value of the shear viscosity is set to $\eta/s=0.08$ and the bulk pressure is neglected. The NEOS-BQS equation of state Monnai:2019hkn is used during the hydrodynamic evolution. When the energy density of the QGP medium drops to the freeze-out energy density 0.4 GeV/fm3, the QGP converts to soft thermal hadrons via the Cooper-Frye prescription Cooper:1974mv and the iSS sampler module. Finally, these soft hadrons are injected into SMASH for further scattering.
For the polarization part, based on the local thermal equilibrium assumption, the polarization vector for spin-1/2 particles can be calculated with the modified Cooper-Frye formula (Becattini:2013fla, ; Fang:2016vpj, ):
$P^{\mu}=\frac{\int d\Sigma^{\alpha}p_{\alpha}f\hat{P}^{\mu}}{\int d\Sigma^{\alpha}p_{\alpha}f}.$ (25)
Here, $f$ is the Fermi-Dirac distribution, and $d\Sigma^{\alpha}$ represents the freeze-out hypersurface elements, determined by the Cornelius routine. The polarization vector $\hat{P}^{\mu}$ can be written as:
$\hat{P}^{\mu}=\hat{P}_{\omega,\mathrm{kin}}^{\mu}+\hat{P}_{\xi,\mathrm{kin}}^{\mu}+\hat{P}_{T}^{\mu}.$ (26)
The first two terms are the velocity-gradient terms in Eq. (14) and Eq. (15). The last term, $\hat{P}_{T}^{\mu}$, represents the polarization caused by the temperature gradient ($T$-gradient) and is given by:
$\hat{P}_{T}^{\mu}=-\frac{1}{2m}\epsilon^{\mu\nu\sigma\tau}(1-f)\omega_{\nu\sigma}^{T}p_{\tau},$ (27)
where the $T$-gradient vorticity is defined as
$\omega_{\mu\nu}^{T}=\frac{1}{2T^{2}}(u^{\nu}\partial^{\mu}T-u^{\mu}\partial^{\nu}T).$ (28)
This term is the sum of the $T$-gradient terms in Eqs.(14) and (15).
There are several differences between the hydrodynamic simulation and the blast wave model. On the one hand, the calculation of the polarization in hydrodynamics does not assume an isothermal condition. Although the temperature on the freeze-out hypersurface is almost constant if we ignore net baryon density, this does not mean the temperature gradient on the hypersurface disappears. For instance, numerical determination of the freeze-out hypersurface in hydrodynamic simulation relies on temperature gradient which provides a normal vector to the freeze-out hypersurface Huovinen:2012is . Therefore, the polarization in hydrodynamic simulation includes the the temperature-gradient contribution. On the other hand, the freeze-out hypersurface in Eq. (25) is not the same as the freeze-out hypersurface in the blast wave model, which only considers time-like terms (such as $\tau dxdyd\eta$) or the iso-proper-time freeze-out. In hydrodynamic simulation, the hypersurface also contains spatial components and depends on the evolution of the QGP medium. Consequently, all types of vorticity in hydrodynamic simulation depend on the evolution time $\tau$, while in the blast wave model, only the vorticity at the final moment is considered.
In order to make a more realistic prediction of the in-plane transverse polarization $P^{x}$, we consider the s-quark equilibrium scenario, where the s-quark mass is set to $m_{s}=$0.5 GeV. We also set the $\hat{t}-$vector in $\hat{P}_{\xi}^{\mu}$ in Eq. (15) as the fluid four-velocity $u^{\mu}$ in our calculation.
Figure 8: The azimuthal angle dependence of the longitudinal polarization $P^{z}$ (left panel) and the out-of-plane polarization $P_{H}=-P^{y}$ (right panel), obtained from hydrodynamic simulation including the kinetic vorticity, shear vorticity, and temperature-gradient contributions in Au+Au collisions at 200 GeV. Experimental data are taken from the STAR Collaboration.
Figure 9: The transverse momentum dependence of the second Fourier coefficient of the longitudinal polarization, $\left<P^{z}\sin(2\phi_{p})\right$> (left panel), and the local transverse polarization $P_{H}$ (right panel), obtained from hydrodynamic simulation with the kinetic vorticity, shear vorticity, and temperature-gradient sources in Au+Au collisions at 200 GeV. The experimental data are taken from STAR collaboration.
Figure 10: The centrality dependence of the second Fourier coefficient of the longitudinal polarization, $\left<P^{z}\sin(2\phi_{p})\right$> (left panel), and the out-of-plane polarization $P_{H}$ (right panel), from hydrodynamic calculation that includes the kinetic vorticity, shear vorticity, and temperature-gradient components in Au+Au collisions at 200 GeV. The experimental data are obtained from the STAR Collaboration.
Figure 11: First panel: hydrodynamic prediction for the in-plane polarization $P^{x}$ as a function of the azimuthal angle; Second and third panels: hydrodynamic prediction for the second Fourier coefficient, $\left<P^{x}\sin(2\phi_{p})\right$>, as functions of $p_{T}$ and centrality. The contributions from the kinetic vorticity, shear stress tensor and temperature-gradient are shown separately.
In Fig. 8, we present the results for the azimuthal angle ($\phi_{p}$) dependence of the longitudinal polarization $P^{z}$ and the out-of-plane polarization $P_{H}$ from different sources. It can be found that the hydrodynamic calculation can also quantitatively describe the $P^{z}$ data, showing similar results to those from previous hydro studies and from the blast wave model. Among different sources, the shear-induced polarization gives the correct $\sin(2\phi_{p})$ behavior compared to the data, and the shear term is the dominant contribution. However, the polarizations by the kinetic vorticity and by the temperature-gradient have the same sign but with different magnitudes. Interestingly, the magnitude of the polarization from the temperature-gradient is even larger than that from the kinetic vorticity. This indicates that the contribution from the temperature gradient cannot be simply neglected, unlike what is assumed in the blast wave model. Note that the magnitude of the data is much smaller than those of the temperature-gradient, kinetic vorticity, and shear contributions, it looks like the small signal is the result of a sum over several large contributions and any small change of one large contribution could flip the sign of the signal.
In Figs. 9 and 10, we present the $p_{T}$ and centrality dependence of the second Fourier coefficient of the longitudinal polarization $\left<P^{z}\sin(2\phi_{p})\right$>, and out-of-plane polarization $P_{H}$ in Au+Au collisions at 200 GeV. The results demonstrate that the hydrodynamic simulation provides a quantitative description of the experimental data. For the longitudinal polarization $\left<P^{z}\sin(2\phi_{p})\right$>, again, the shear contribution gives the correct slope in $p_{T}$ and centrality spectra, while the temperature-gradient contribution gives an opposite slope, although they have almost equal magnitude: two contributions are in competition. For the out-of-plane polarization $P_{H}$, the hydrodynamic calculation does not produce its $\phi_{p}$ dependence observed in experiments as shown in Fig. 8. However, the magnitude of the total $P_{H}$ can be described by the hydrodynamic model, as shown in Fig. 9 and Fig. 10. It can be found that $P_{H}$ mainly comes from the kinetic vorticity contribution, which originates from the initial angular momentum.
The numerical results in Fig. 8-10 show that our hydrodynamic model can provide a description of the polarization data and capture the gross features of the fireball’s dynamical evolution. Then we try to make a prediction for the in-plane polarization $P^{x}$ as functions of $\phi_{p}$, $p_{T}$ and centrality in Fig. 11. Compared to the prediction by the blast wave model, the hydrodynamic simulation result of $P^{x}$ shows an overall opposite sign. This difference mainly arises from a significant contribution from the temperature-gradient component in $P^{x}$, while the kinetic and shear terms show the same sign as the blast wave prediction. Similar to $P^{z}$, the magnitude of $P^{x}$ in total is much smaller than those of the temperature-gradient and shear contributions, it looks like the small signal is the result of the difference between two large contributions. The sign of $P^{x}$ can be flipped by any small change of one large contributions. It has been shown in (Lin:2024svh, ; Li:2025zbj, ) that, in an interacting system, the contribution of the velocity gradient and temperature gradient will be modified, which may enhance or diminish the result after cancellation. The role played by the temperature-gradient in the in-plane polarization will have to be elucidated through further experimental measurements.
VII Summary
We performed a comprehensive analysis of in-plane transverse polarization in heavy-ion collisions, offering both analytical solutions and numerical results. For the in-plane polarization, the hydrodynamic simulation predicts an opposite sign compared to the blast wave model prediction, mainly because of the contribution from the temperature-gradient component. In the hydrodynamic simulation, similar to the longitudinal polarization, the magnitude of the total result for the in-plane polarization is much smaller than those of the temperature-gradient and shear contributions, i.e. the small signal for the in-plane polarization is the result of the difference between two large contributions. The sign of the signal can be easily flipped under any small change of either large contribution. Future experiments may help to determine the impact of the temperature-gradient on the in-plane polarization.
Acknowledgements.
A large part of this work was completed during the time when Q.W. was a visiting professor in the nuclear theory group at McGill University. The authors thank X.-L. Sheng and S. Lin for helpful discussion. This work is supported in part by the National Natural Science Foundation of China (NSFC) under Grant No. 12135011, and in part by the Natural Sciences and Engineering Research Council of Canada (NSERC) [SAPIN-2020-00048, SAPIN-2024-00026].